OpenStudy (anonymous):

if log2(x)=log2(a)+log2(b)-log2(c) then what is x?

OpenStudy (jack1):

log2(x)=log2(a)+log2(b)-log2(c) log2(x)=log2 [(a*b)/c] so x = (a * b / c)

OpenStudy (caozeyuan):

the upstairs have made a perfect solution

OpenStudy (jhannybean):

lol..

OpenStudy (caozeyuan):

it's chinglish really, in CHInese forum we oftener use the word 楼上,which means "upstairs" to refer to people who appeared before you

OpenStudy (unklerhaukus):

\[\large\boxed{\log_b p+\log_bq=\log_bpq}\\\large\boxed{\log_b p-\log_bq=\log_b\frac pq}\]